OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (1, 3) and (3, 7)? y = −2x + 1 y = −2x − 1 y = 2x + 1 y = 2x − 1

OpenStudy (eujc21):

m = (y2-y1)/(x2-x1)

OpenStudy (eujc21):

y - y1 = m(x - x1)